An aqueous solution of phosphoric acid, that is 0.500 percent by weight phosphoric acid, has a density of 1.0010 g/mL. A student determines that the freezing point of this solution is -0.129 °C.

Based on the observed freezing point, what is the percent ionization of the acid and the value of Ka1? Assume that only the first ionization of the acid is important.
Kf for H2O is 1.86 °C/m.

% ionized =    %
Ka1   =   
Answer #1

Answer:-

Given:-

freezing point of pure water (H2O) Tf0 = 0 0C

freezing point of solution Tf = - 0.129 0C

decrease in freezing point  Tf  = Tf0 - Tf = 0 - (- 0.129) = 0.129 0C

freezing point depression constant of water (H2O) (Kf) = 1.86 0C /m

density of phosphoric solution = 1.0010 g / ml

wt. of phosphoric acid (H3PO4) by (w/v) = 0.500

molar mass of phosphoric acid (H3PO4) = 98 g mol-1

As we know that 0.500 percent by weight of phosphoric acid (H3PO4) means that 0.500 g of phosphoric acid (H3PO4) is present in 100 g of aqueous solution.

the wt. of phosphoric acid (H3PO4) = 0.500 g

wt of water (H2O) = 100 - 0.500 = 99.5 g = 0.0995 Kg

therefore

no. of moles of phosphoric acid (H3PO4) (n) = wt. of phosphoric acid (H3PO4) / molar mass of phosphoric acid (H3PO4)

no. of moles of phosphoric acid (H3PO4) (n) = 0.500 g / 98 g mol-1

no. of moles of phosphoric acid (H3PO4) (n) = 0.0051 mol

As we know that

molaity of phosphoric acid (H3PO4) solution (m) = no. of moles of phosphoric acid (H3PO4) (n) / wt of water (H2O) in Kg

molaity of phosphoric acid (H3PO4) solution (m) = 0.0051 mol / 0.0995 Kg

molaity of phosphoric acid (H3PO4) solution (m) = 0.05126 m

According to the formula

Tf  = i Kf m

0.129 0C = i 1.86 0C /m 0.05126 m

Van't Hoff factor (i) =  0.129 0C / 1.86 0C /m 0.05126 m

Van't Hoff factor (i) =   0.129 / 0.0953436

Van't Hoff factor (i) = 1.353 ----------------------------------------------- (1)

As we know that

mass of solution = volume of solution density of solution

volume of solution phosphoric acid (H3PO4) = mass of (H3PO4) solution /  density of (H3PO4) solution

volume of solution phosphoric acid (H3PO4) = 100 g / 1.0010 g / ml

volume of solution phosphoric acid (H3PO4) = 99.9 ml = 0.0999 L

Also we know that

Molarity of phosphoric acid (H3PO4) = no. of moles of (H3PO4) (n) / volume of solution (H3PO4) in Liter

Molarity of phosphoric acid (H3PO4) =  0.0051 mol / 0.0999 L

Molarity of phosphoric acid (H3PO4) = 0.0512 M

As we know that phosphoric acid (H3PO4) ionized as follows:-

   H3PO4       H+ + H2PO4-

Initial 0.0512 M 0 0   

Change - + +

Equilibrium (0.0512 M - )    + +

since is the degree of dissociation of phosphoric acid (H3PO4).

therefore

total moles of equilibrium = (0.0512 M - ) + + = 0.0512 M +

Also we know that

Van't Hoff factor (i) = 0.0512 M + /  0.0512 M ---------------------------------- (2)

from equation no. 1 and 2 we get

1.353 = 0.0512 M + /  0.0512 M

0.0512 M +   = 0.0512 M    1.353

0.0512 M +   = 0.0692736 M

= 0.0692736 M - 0.0512 M

= 0.01807

therefore

percent ionization of  phosphoric acid (H3PO4) (%) = degree of dissociation of phosphoric acid ()   100

percent ionization of  phosphoric acid (H3PO4) (%) = 0.01807 ​​​​​​​ 100

percent ionization of  phosphoric acid (H3PO4) (%) = 1.807 % ( i.e the answer)

As we know that

H3PO4    H+ + H2PO4-

Initial 0.0512 M 0 0   

Change - 0.01807 M +0.01807 M +0.01807 M

Equilibrium (0.0512 M - 0.01807 M )    +0.01807 M +0.01807 M

0.03313 M +0.01807 M +0.01807 M
therefore we get

molar concentration of [H3PO4]= 0.03313 M

molar concentration of [H+ ] = 0.01807 M

molar concentration of [H2PO4-] = 0.01807 M

As we know that

equilibrium constant of H3PO4 (K1) = [H+ ][H2PO4-] / [H3PO4]

equilibrium constant of H3PO4 (K1) =  0.01807   0.01807 / 0.03313  

equilibrium constant of H3PO4 (K1) =  0.0003265249 / 0.03313  

equilibrium constant of H3PO4 (K1) = 0.009856

equilibrium constant of H3PO4 (K1) = 9.856 10-3 ( i.e the answer)